Use Lagrange multipliers to find the maximum and minimum values of f(x,y)=x−2y subject to the constraint x^2+y^2=5, if such values exist.

maximum = ?

minimum = ?
1
Expert's answer
2020-04-07T13:23:57-0400

Given "f(x,y) = x - 2y" .

Let "g(x,y) = x^{2} + y^{2} - 5" and

let "F = f(x,y)+\lambda g(x,y)" , where "\lambda" is the Lagrangian multiplier.

"F = x - 2y + \lambda(x^{2} + y^{2} -5)"


To find the maximum and minimum values, we have to solve the system

"\dfrac{\partial F}{\partial x} = 0","\dfrac{\partial F}{\partial y} = 0" and "\dfrac{\partial F}{\partial \lambda} = 0" .


"1 + 2x\lambda = 0\\
-2 + 2y\lambda = 0\\
x^{2} + y^{2} = 5."


From the first and the second equation we get,

"x = -\dfrac{1}{2\lambda}\\

y = \dfrac{1}{\lambda}\\"


Using these values in the third equation we get,

"\dfrac{1}{4\lambda^{2}} + \dfrac{1}{\lambda^{2}} = 5\\

\lambda^{2} = \dfrac{1}{4}\\

\lambda = \pm\dfrac{1}{2}\\"


When "\lambda = -\dfrac{1}{2}" ,

"x = 1" and "y = -2"

When "\lambda = \dfrac{1}{2}" ,

"x = -1" and "y = 2"


The maximum and minimum values of f(x,y) are respectively,

"f(1,-2) = 1+4 = 5\\

f(-1,2) = -1-4 = -5\\"